OpenStudy (anonymous):

The grades on the last history exam had a mean of 80%. Assume the population of grades on history exams is known to be normally distributed with a standard deviation of 5%. What percent of students earn a score between 70% and 80%?

OpenStudy (anonymous):

A. 0.022 B. 0.477 C. 0.50 D. 0.522

OpenStudy (anonymous):

Please help!!

jimthompson5910 (jim_thompson5910):

do you have a TI-83 or TI-84 calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Can you walk me through it?

jimthompson5910 (jim_thompson5910):

ok let me find one

jimthompson5910 (jim_thompson5910):

ok so here is one calculator you can use http://www.mathportal.org/calculators/statistics-calculator/normal-distribution-calculator.php you type 80 in the \(\Large \mu\) box for the mean type 5 for the \(\Large \sigma\) box (standard deviation) then you pick the option that has \(\Large P( \square <X < \square ) \) typing 70 and 80 in the two boxes (smaller number goes in the left box)

jimthompson5910 (jim_thompson5910):

that calculator is really cool because you can click the checkbox "show explanation" and it gives you a nice visual step by step

OpenStudy (anonymous):

Ahhh! Thank you so much! I had the formula, put didn't know what the symbols meant.

jimthompson5910 (jim_thompson5910):

yeah it's probably a bit odd to see greek symbols being used \[\Large \mu = \text{mu}\] the "mu" is basically the greek letter m (well it's like the letter m in the latin alpabet) think M for Mean \[\Large \sigma = \text{sigma}\] "sigma" is closely linked to the letter S S for Standard deviation

jimthompson5910 (jim_thompson5910):

mu and sigma here are the lower case greek forms

OpenStudy (anonymous):

Ok, can you help with another similar to this one?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

In a study of 250 adults, the mean heart rate was 70 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 12 beats per minute. What is the 99% confidence interval for the mean beats per minute? A. 68.9-76.3 B. 70-72 C. 61.2-72.8 D. 68-72

jimthompson5910 (jim_thompson5910):

what is the z critical value for the 99% confidence interval ?

OpenStudy (anonymous):

2.576 I think

jimthompson5910 (jim_thompson5910):

very good, since the solution to P( -k < X < k ) = 0.99 is roughly k = 2.576

jimthompson5910 (jim_thompson5910):

the mean heart rate was 70 beats per minute so xbar = 70

jimthompson5910 (jim_thompson5910):

"Assume the population of heart rates is known to be approximately normal with a standard deviation of 12 beats per minute" sigma = 12

jimthompson5910 (jim_thompson5910):

"In a study of 250 adults" ---> n = 250

jimthompson5910 (jim_thompson5910):

You will now use the formulas \[\Large L = \overline{ x} - z^{*} \frac{\sigma}{\sqrt{n}}\] \[\Large U = \overline{ x} + z^{*} \frac{\sigma}{\sqrt{n}}\]

jimthompson5910 (jim_thompson5910):

\[\Large \overline{x} = 70\] \[\Large \sigma = 12\] \[\Large z^{*} \approx 2.576\] \[\Large n = 250\]

jimthompson5910 (jim_thompson5910):

to get the confidence interval (L,U)

OpenStudy (anonymous):

I got D, 68-72, is that right?

jimthompson5910 (jim_thompson5910):

\[\Large L = \overline{ x} - z^{*} \frac{\sigma}{\sqrt{n}}\] \[\Large L = 70 - 2.576* \frac{12}{\sqrt{250}}\] \[\Large L = 68.04495\] \[\Large L = 68\] -------------------------------------------------- \[\Large U = \overline{ x} + z^{*} \frac{\sigma}{\sqrt{n}}\] \[\Large U = 70 + 2.576* \frac{12}{\sqrt{250}}\] \[\Large U = 71.95505\] \[\Large U = 72\] -------------------------------------------------- I agree, the confidence interval is roughly (L,U) = (68,72)

OpenStudy (anonymous):

Thank you!! You have helped me so much!

jimthompson5910 (jim_thompson5910):

I'm glad I could help out